A basket ball team has won 30 games of 40 played.How many of the remaining 30 games must it win to have a 80% win record for the season ?
Answer by RAY100(1637)

Easiest to make a quick TABLE
'
won,,,,,,,,,,,,total,,,,,,% win
.
30,,,,,,,,,,,,,40 ,,,,,,,,,,(75 %)
.
x,,,,,,,,,,,,,,30,,,,,,,,,,,,,?
.
30+x,,,,,,,,,,,70,,,,,,,,,,,,80%(goal),,,,,,,TOTAL
.
But 80 % of 70 =(.8*70)= 56 =(30+x),,,,,x=26 more games to win to get 80% TOTAL
.